# A modern perspective on rational homotopy theory

### v1

@@ 4.4. Fibrations of $1$-reduced simplicial sets
We prove a characterization that detects when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration and helps us disprove the converse of Proposition 4.12. Our result is the $1$-reduced analog of [GJ99, Lemma 6.6]. Recall from Lemma 3.13 that, for $0\leq k\leq 2$, the map ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks the unique 00-simplex of the simplicial $2$-sphere.
@@
Let $f\colon X\rightarrow Y$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For $0\leq k\leq 2$, let $j_{2,k}\colon\Lambda[2,k]\rightarrow\Delta[2]$ denote the horn inclusion. The following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is a Kan fibration.
  2. (2)
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
@@
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, a lifting problem of the type
@@
(II) ${{\Lambda[2,k]}}$${X}$${{\Delta[2]}}$${{Y.}}$$\scriptstyle{j_{2,k}}$$\scriptstyle{f}$
If $f$ is a Kan fibration, then every lifting problem of type (II) has a solution. Thus, every lifting problem of type (I) has a solution, and $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
Conversely, if $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$, then every lifting problem of type (II) has a solution. If $n\neq 2$ and $0\leq t\leq n$, then solving a lifting problem
${{\Lambda[n,t]}}$${X}$${{\Delta[n]}}$${{Y}}$$\scriptstyle{j_{n,t}}$$\scriptstyle{f}$
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to solving a lifting problem
${{{R}_{1}\left(\Lambda[n,t]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right)}}$${{Y.}}$$\scriptstyle{{R}_{1}\left(j_{n,t}\right)}$$\scriptstyle{f}$
By Lemma 4.8, ${R}_{1}\left(j_{n,t}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $n\neq 2$, so our lifting problem has a solution because $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Overall, we conclude that $f$ is a Kan fibration. ∎
@@
Let $B$ be a $1$-reduced simplicial set. Then, a map $h\colon B\rightarrow S_{\mathrm{simp}}^{2}$ is constant if and only if
$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
is not an epimorphism.
@@
If $h$ is constant, then so is $h_{\ast}$. If $h$ is not constant, then there is a $2$-simplex $\omega$ of $B$ such that $h(\omega)$ is the unique non-degenerate $2$-simplex $\epsilon$ of $S_{\mathrm{simp}}^{2}$. As $B$ is $1$-reduced, Corollary 2.20 allows us to view $\omega$ as a map
$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow B.$
Then, we compute that
$h_{\ast}\left([\omega]\right)=[h(\omega)]=[\epsilon],$
which is a generator of $\pi_{2}(S_{\mathrm{simp}}^{2})\cong\mathbb{Z}$. Thus, $h_{\ast}$ is an epimorphism. ∎
@@
For $0\leq k\leq 2$, the map of $1$-reduced simplicial sets ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ but not a Kan fibration.
@@
We first show that ${R}_{1}\left(j_{2,k}\right)$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For an acyclic cofibration $\gamma\colon A\rightarrow B$ in ${\mathcal{SS}\mathrm{ets}}_{1}$ and a lifting problem
${A}$${{\Delta[0]}}$${B}$${{S_{\mathrm{simp}}^{2},}}$$\scriptstyle{c}$$\scriptstyle{\gamma}$$\scriptstyle{{R}_{1}\left(j_{2,k}\right)}$$\scriptstyle{g}$$\scriptstyle{h}$
we claim that the unique option for $g$ solves our lifting problem. The upper left triangle has to commute. To show that the lower right triangle also commutes, we show that $h$ is constant. By Lemma 4.14, we show that
$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
is not an epimorphism; in fact, we show that $h_{\ast}$ is constant. Firstly, as $c_{\ast}$ is constant, so is
$h_{\ast}\gamma_{\ast}=(h\gamma)_{\ast}=({{R}_{1}\left(j_{2,k}\right)}c)_{\ast}={{R}_{1}\left(j_{2,k}\right)}_{\ast}c_{\ast}.$
@@
Now, we apply the characterization from Proposition 4.13 to argue that ${R}_{1}\left(j_{2,k}\right)$ is not a Kan fibration because it does not have the right lifting property with respect to itself. Indeed, the lifting problem
[…]
@@
In the presence of a fibrancy condition on the target, Proposition 4.13 can be strengthened. Namely, we establish the following characterization result, which is the $1$-reduced analog of [GJ99, Corollary 6.9].
@@
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
@@
In our proof, we use the construction of the homotopy groups of Kan complexes from [GJ99, Chapter I.7].
@@
Now, Proposition 4.13 says that (1) and (2) are equivalent. We show that (2) and (3) are equivalent. Firstly, suppose that $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$. As $Y$ is a Kan complex, every element of $\pi_{2}(Y)$ is of the form $[\omega]$ for a map
$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow Y.$
We solve the lifting problem
[…]
to produce an element $[\alpha]\in\pi_{2}(X)$ such that
$f_{\ast}([\alpha])=[f\alpha]=[\omega]\in\pi_{2}(Y).$
@@
where
$f_{\ast}([\beta])=[f\beta]=[\eta]\in\pi_{2}(Y)$
by means of the pointed homotopy
$H\colon S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}\rightarrow Y$
starting from $Hd^{1}=f\beta$ and ending at $Hd^{0}=\eta$. By Lemma 2.15, $S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}$ is $1$-reduced. As $d^{1}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, there exists a lift
@@
Then, $\psi=gd^{0}$ solves our lifting problem because
$f\psi=fgd^{0}=Hd^{0}=\eta.$
The fibrancy condition in Proposition 4.16 is the reason why we could not have used it for our example of a pathological fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Proposition 4.15. Indeed, we conclude this section by showing that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex. To that end, we need a counting lemma.
@@
Hence, in either case, $s_{k}(x)$ has at most $2$ non-degenerate faces.
@@
For $10\leq i\leq n+$ 1, let $d_{i}s_{k}(x)$ be the $i$-th face of $s_{k}(x)$. The result follows from the simplicial identities:
@@
Write $\epsilon$ for the unique non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. For $10\leq k\leq n+$ 1, consider the lifting problem
@@
where $h$ is the horn that picks out the $n$-simplices $\left(\epsilon,\dots,\epsilon\right)$. Any simplex $g$ that fills in this horn would be a degenerate $(n+1)$-simplex with $n+1$ non-degenerate faces, which is impossible for $n\geq 2$ by Lemma 4.17. ∎
@@ 5. The simplicial model structure on $1$-reduced simplicial sets
A convenient feature of the model categories ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is that they are simplicial, so they carry more structure for us to use. This feature motivates us to endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure. We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors of $1$-reduced simplicial sets. Then, we verify that our definitions endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
@@ 5.1. Simplicial model categories
A reference on simplicial model categories is [Hir03, Chapter 9].
Definition 5.1 ([Hir03, Definition 9.1.2]).
Let $\mathcal{M}$ be a model category that is enriched over simplicial sets; in particular, for two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)$. Then, $\mathcal{M}$ is a simplicial model category if $\mathcal{M}$ satisfies the following axioms.
@@
    (Adjointness axiom) For two objects $X$ and $Y$ of $\mathcal{M}$ and a simplicial set $Z$, there is a tensor object $X\otimes Z$ of $\mathcal{M}$ and a cotensor object $Y^{Z}$ of $\mathcal{M}$ such that there are isomorphisms of simplicial sets
    $\operatorname{Map}_{\mathcal{M}}\left(X\otimes Z,Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)\cong\operatorname{Map}_{\mathcal{M}}\left(X,Y^{Z}\right)$
    that are natural in $X$, $Y$, and $Z$. Note that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)$ is a function complex of simplicial sets; see Example 5.2.
@@
    (Homotopy lifting-extension axiom) If $i\colon A\rightarrow B$ is a cofibration in $\mathcal{M}$ and $p\colon X\rightarrow Y$ is a fibration in $\mathcal{M}$, then the induced map
    $\operatorname{Map}_{\mathcal{M}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{\mathcal{M}}\left(A,X\right)\underset{\operatorname{Map}_{\mathcal{M}}\left(A,Y\right)}{\times}\operatorname{Map}_{\mathcal{M}}\left(B,Y\right)$
    is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak equivalence in $\mathcal{M}$.
Example 5.2 ([Hir03, Example 9.1.13]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(X\times\Delta[n],Y\right)$;
@@
    the cotensor of $X$ with $Z$ to be the simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,X\right)$.
Example 5.3 ([Hir03, Example 9.1.14]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,Y\right)$ as the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},Y\right)$;
@@
    the tensor of $X$ with $Z$ to be the pointed simplicial set $X\times Z_{+}$; and
@@
    the cotensor of $X$ with $Z$ to be the pointed simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+},X\right)$.
5.2. Function complexes, tensors, and cotensors of $1$-reduced simplicial sets
We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors in ${\mathcal{SS}\mathrm{ets}}_{1}$. We define

### v2

@@ 4.4. Pathological fibrations of $1$-reduced simplicial sets
We prove a characterization that detects when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration and helps us disprove the converse of Proposition 4.11; our result is the $1$-reduced analog of [23, Lemma 6.6]. Recall from Lemma 3.13 that, for $0\leq k\leq 2$, the map ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks the unique 0-simplex of the simplicial $2$-sphere.
@@
Let $f$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For $0\leq k\leq 2$, let $j_{2,k}\colon\Lambda[2,k]\rightarrow\Delta[2]$ denote the horn inclusion. Then, $f$ is a Kan fibration if and only if $f$ has the right lifting property against ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
@@
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, a lifting problem of the type
@@
[…]
If $f$ is a Kan fibration, then every lifting problem of type (II) has a solution. Thus, every lifting problem of type (I) has a solution, and $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
Conversely, if $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$, then every lifting problem of type (II) has a solution. If $n\neq 2$ and $0\leq t\leq n$, then solving a lifting problem
[…]
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, to solving a lifting problem
[…]
By Lemma 4.7, ${R}_{1}(j_{n,t})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $n\neq 2$, so our lifting problem has a solution because $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Overall, we conclude that $f$ is a Kan fibration. ∎
@@
Let $B$ be a $1$-reduced simplicial set. Then, a map $h\colon B\rightarrow S_{\mathrm{simp}}^{2}$ is constant if and only if $h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$ is not an epimorphism.
@@
If $h$ is constant, then so is $h_{\ast}$. If $h$ is not constant, then there is a $2$-simplex $\omega$ of $B$ such that $h(\omega)$ is the unique non-degenerate $2$-simplex $\epsilon$ of $S_{\mathrm{simp}}^{2}$. As $B$ is $1$-reduced, Corollary 2.16 allows us to view $\omega$ as a map $\omega\colon S_{\mathrm{simp}}^{2}\rightarrow B$. Then, we compute that $h_{\ast}([\omega])=[h(\omega)]=[\epsilon]$, which is a generator of $\pi_{2}(S_{\mathrm{simp}}^{2})\cong\mathbb{Z}$. ∎
@@
For $0\leq k\leq 2$, the map of $1$-reduced simplicial sets ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ but not a Kan fibration.
@@
We first show that ${R}_{1}(j_{2,k})$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For an acyclic cofibration $\gamma\colon A\rightarrow B$ in ${\mathcal{SS}\mathrm{ets}}_{1}$ and a lifting problem
[…]
we claim that the unique option for $g$ solves our lifting problem. The upper left triangle has to commute. To show that the lower right triangle also commutes, we show that $h$ is constant. By Lemma 4.13, we show that $h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$ is not an epimorphism; in fact, we show that $h_{\ast}$ is constant. As $c_{\ast}$ is constant, so is
$h_{\ast}\gamma_{\ast}=(h\gamma)_{\ast}=({{R}_{1}(j_{2,k})}c)_{\ast}={{R}_{1}(j_{2,k})}_{\ast}c_{\ast}.$
@@
Now, we apply the characterization from Proposition 4.12 to argue that ${R}_{1}(j_{2,k})$ is not a Kan fibration because it does not have the right lifting property with respect to itself. Indeed, the lifting problem
$\Delta[0]$$\Delta[0]$$S_{\mathrm{simp}}^{2}$$S_{\mathrm{simp}}^{2}$$\scriptstyle{{R}_{1}(j_{2,k})}$$\scriptstyle{{R}_{1}(j_{2,k})}$$\scriptstyle{g}$
@@
In the presence of a fibrancy condition on the target, Proposition 4.12 can be strengthened. Namely, we establish the following characterization result, which is the $1$-reduced analog of [23, Corollary 6.9].
@@
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
@@
The unique map $X\rightarrow\Delta[0]$ factors in ${\mathcal{SS}\mathrm{ets}}_{1}$ as the composite of fibrations $X\xrightarrow{f}Y\rightarrow\Delta[0]$. Hence, $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$, thus a Kan complex by the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 3.17. Now, Proposition 4.12 says that (1) and (2) are equivalent. We show that (2) and (3) are equivalent. Firstly, suppose that $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$. As $Y$ is a Kan complex, every element of $\pi_{2}(Y)$ is of the form $\omega]$ for a map $\omega\colon S_{\mathrm{simp}}^{2}\rightarrow Y$. We solve the lifting problem
@@
to produce an element $\alpha]\in\pi_{2}(X)$ such that $f_{\ast}([\alpha])=[f\alpha]=[\omega]$.
@@
where $f_{\ast}([\beta])=[f\beta]=[\eta]$ by means of the homotopy $H\colon S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}\rightarrow Y$ starting from $Hd^{1}=f\beta$ and ending at $Hd^{0}=\eta$. By Lemma 2.11, $S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}$ is $1$-reduced. As $d^{1}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, there exists a lift
@@
Then, $\psi=gd^{0}$ solves our lifting problem because $f\psi=fgd^{0}=Hd^{0}=\eta$. ∎
The fibrancy condition in Proposition 4.15 is the reason why we could not have used it to streamline our proof of Proposition 4.14. Indeed, we conclude this section by showing that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex. To that end, we need a counting lemma.
@@
Thus, the degenerate $(n+1)$-simplex $s_{k}(x)$ has at most $2$ non-degenerate faces.
@@
For $0\leq i\leq n+1$, let $d_{i}s_{k}(x)$ be the $i$-th face of $s_{k}(x)$. The result follows from the simplicial identities
@@
Our counting lemma allows us to prove that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex.
@@
Write $\epsilon$ for the unique non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. For $0\leq k\leq n+1$, consider the lifting problem
@@
where $h$ is the horn that picks out the $n$-simplices $(\epsilon,\dots,\epsilon)$. Any simplex $g$ that fills in this horn would be a degenerate $(n+1)$-simplex with $n+1$ non-degenerate faces, which is impossible for $n\geq 2$ by Lemma 4.16. ∎
@@ 5. The simplicial model structure on $1$-reduced simplicial sets
A convenient feature of the model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ is that they are simplicial, so they carry more structure for us to use. In particular, between any two objects, we have an explicit construction of a function complex, which facilitates the study of left Bousfield localizations of these model categories.
In this section, we endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure. This model structure makes our subsequent left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ more convenient in practice. Moreover, as left Bousfield localizations preserve the simplicial structure of a model category, the entire family of model structures on $1$-reduced simplicial sets inherits the simplicial structure in this section, so we get significant traction out of our argument.
@@ 5.1. Simplicial model categories
A reference on simplicial model categories is [26, Chapter 9]. We reformulate axiom (2) in the definition from [26, Definition 9.1.2] for our purposes using [26, Proposition 9.3.7].
Definition 5.1.
Let $\mathcal{M}$ be a model category that is enriched in simplicial sets; in particular, for two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$. Then, $\mathcal{M}$ is a simplicial model category if $\mathcal{M}$ satisfies the following axioms.
@@
    (Adjointness axiom) For two objects $X$ and $Y$ of $\mathcal{M}$ and a simplicial set $Z$, there is a tensor object $X\otimes Z$ of $\mathcal{M}$ and a cotensor object $Y^{Z}$ of $\mathcal{M}$ satisfying the natural isomorphisms of simplicial sets
    $\operatorname{Map}_{\mathcal{M}}(X\otimes Z,Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{\mathcal{M}}(X,Y))\cong\operatorname{Map}_{\mathcal{M}}(X,Y^{Z}).$
    Note that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{\mathcal{M}}(X,Y))$ is a function complex of simplicial sets; see Example 5.2.
@@
    (Homotopy lifting-extension axiom) If $i\colon A\rightarrow B$ is a cofibration in $\mathcal{M}$ and $j\colon L\rightarrow K$ is a cofibration of simplicial sets, then the induced morphism
    $A\otimes K\underset{A\otimes L}{\coprod}B\otimes L\rightarrow B\otimes K$
    is a cofibration in $\mathcal{M}$ that is acyclic if $i$ or $j$ is acyclic.
We present the simplicial structure of the model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$.
Example 5.2 ([26, Example 9.1.13]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(X,Y)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}(X\times\Delta[n],Y)$;
@@
    the cotensor of $X$ with $Z$ to be the simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,X)$.
Example 5.3 ([26, Example 9.1.14]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$ as the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},Y)$;
@@
    the tensor of $X$ with $Z$ to be the pointed simplicial set $X\wedge Z_{+}$; and
@@
    the cotensor of $X$ with $Z$ to be the pointed simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)$.
5.2. The simplicial model structure
As our left transfer reveals to us, the homotopy theory of $1$-reduced simplicial sets is closer to that of pointed simplicial sets. Thus, it is the simplicial structure from Example 5.3 that we modify to get a simplicial structure on ${\mathcal{SS}\mathrm{ets}}_{1}$. In fact, our modification guarantees that the homotopy lifting-extension axiom for ${\mathcal{SS}\mathrm{ets}}_{1}$ is identical to that for $\mathcal{SS}\mathrm{ets}$$_{\ast}$, so we circumvent what is often the most difficult part of endowing a model category with a simplicial structure.
For the simplicial structure on the model category ${\mathcal{SS}\mathrm{ets}}_{1}$, we define